# Dynamic Learning Rate for Deep Reinforcement Learning: A Bandit Approach

### v1

@@ Abstract
In Deep Reinforcement Learning models trained using gradient-based techniques, the choice of optimizer and its learning rate are crucial to achieving good performance: higher learning rates can prevent the model from learning effectively, while lower ones might slow convergence. Additionally, due to the non-stationarity of the objective function, the best-performing learning rate can change over the training steps. To adapt the learning rate, a standard technique consists of using decay schedulers. However, these schedulers assume that the model is progressively approaching convergence, which may not always be true, leading to delayed or premature adjustments. In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a meta-learning approach that selects the learning rate based on the agent’s performance during training. LRRL is based on a multi-armed bandit algorithm, where each arm represents a different learning rate, and the bandit feedback is provided by the cumulative returns of the RL policy to update the arms’ probability distribution. Our empirical results demonstrate that LRRL can substantially improve the performance of deep RL algorithms.
@@ 1 Introduction
Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.33, 19, 39, 40). While ANNs were previously used as value function approximators [29], the introduction of Deep Q-Networks (DQN) by [24, 25] marked a significant breakthrough by improving learning stability through two mechanisms: the target network and experience replay.
@@
The experience replay (see 22) stores the agent’s interactions within the environment, allowing sampling of past interactions in a random way that disrupts their correlation. The target network further stabilizes the learning process by periodically copying the parameters of the learning network. This strategy is crucial because the Bellman update —using estimations to update other estimations— would otherwise occur using the same network, potentially causing divergence. By leveraging the target network, gradient steps are directed towards a periodically fixed target, ensuring more stability in the learning process. Additionally, the learning rate hyperparameter controls the magnitude of these gradient steps in optimizers such as the stochastic gradient descent algorithm, affecting the training convergence.
@@
The learning rate is one of the most important hyperparameters, with previous work demonstrating that decreasing its value during policy finetuning can enhance performance by up to 25% in vanilla DQN [3]. Determining the appropriate learning rate [^1] The terms “learning rate” and “step-size” are often used interchangeably in the literature and they technically refer to the same concept. is essential for achieving good model performance: higher values can prevent the agent from learning, while lower values can lead to slow convergence (see 14, 7, 42). However, finding a learning rate value that improves the model performance requires extensive and computationally expensive testing. In order to adapt its initial choice during training, optimizers such as Adam [17] and RMSProp [37] employ an internal scheme that dynamically adjusts the learning rate, considering, for instance, past gradient information. Nevertheless, various learning rate scheduling strategies can be combined with the optimizer to decrease the learning rate and improve the convergence over the training steps.
@@
Standard learning rate schedulers typically decrease the learning rate based on training progress using, e.g., linear or exponential decay strategies [31, 42]. In the context of RL, this approach can lead to premature or delayed learning rate adjustments, which may hinder the agent’s ability to learn. Unlike supervised learning, RL usually involves generating data by trading off between exploration (discovery of new states) and exploitation (refining of the agent’s knowledge). As the policy improves, the data distribution encountered by the agent becomes more concentrated, but this evolution occurs at a different pace than the overall training progress. For instance, some environments require extensive exploration due to the sparseness of rewards, while others need more exploitation to refine the policy to the complexity of the task. Consequently, a more sophisticated decaying learning rate strategy that accounts for policy performance rather than training steps can significantly enhance learning in deep RL.
@@
In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a method to select the learning rate on the fly for deep RL. Our approach acknowledges that different learning phases require different learning rates, and as such, instead of scheduling the learning rate decay using some blanket approach, we dynamically choose the learning rate using a Multi-Armed Bandit (MAB) algorithm, which accounts for the current policy’s performance. Our method has the advantage of being algorithm-agnostic and applicable to any optimizer, although the results show that it works best when coupled with Adam. We conduct experiments on our approach using baselines provided in the Dopamine framework (see 9). Our results focus on exploiting different settings for LRRL to illustrate its robustness under many possible configurations. Our main contributions are the following:
  • We introduce LRRL, the first approach, to our knowledge, that leverages a multi-armed bandit algorithm to select the learning rate dynamically in deep RL. Our results demonstrate that LRRL achieves competitive performance with or superior to standard deep RL algorithms using fixed baselines or traditional learning rate schedulers.
  • Our results show that LRRL significantly reduces the need for hyperparameter optimization by dynamically selecting from a set of possible learning rates using a multi-armed bandit approach. This method mitigates the need for exhaustive techniques like grid search, as it efficiently adapts the learning rate during training in a single run, rather than requiring multiple runs to test each learning rate individually.
  • We assess the robustness of our method by employing the Adam and RMSProp optimizers with different sets of arms. We also compare the results using stochastic and adversarial multi-armed bandit algorithms in Appendix A.2.
2 Related Work
Multi-armed bandit for (hyper)parameter selection.
Deep RL is known to be overly optimistic in the face of uncertainty [27], and many works have addressed this issue by proposing conservative policy updates [38, 13, 2]. However, when the agent is able to interact with the environment, this optimism can encourage exploration, potentially leading to the discovery of higher returns. Building on this idea, [26] use an adversarial MAB algorithm to trade-off between pessimistic and optimistic policy updates based on the agent’s performance over the learning process. In order to select the learning rate for stochastic gradient MCMC, [10] employ an algorithm based on Successive Halving [16, 15], a MAB strategy that promotes promising arms and prunes suboptimal ones over time. In the context of hyperparameter optimization, Successive Halving has also been used in combination with infinite-arm bandits to select hyperparameters for supervised learning [21, 32]. A key difference between these approaches to hyperparameter optimization for supervised learning and our work is that we focus on selecting the best learning rate from a predefined set, rather than performing an extensive and computationally expensive search over the hyperparameter space. Close to our work, [23] propose Adam with Bandit Sampling (ADAMBS), which employs the Exponential-weight algorithm for Exploration and Exploitation [6] to enhance sample efficiency by incorporating importance sampling within the Adam optimizer. While ADAMBS prioritizes informative samples, our method focuses on RL tasks by dynamically adjusting the learning rate, accelerating learning when the policy is far from optimal, and slowing it down as it converges.
Learning rate adapters/schedulers.
Optimizers such as RMSProp [37] have an adaptive mechanism to update a set of parameters $\theta$ by normalizing past gradients, while Adam [17] also incorporates momentum to smooth gradient steps. However, despite their widespread adoption, these algorithms have inherent limitations in non-stationary environments since they do not adapt to changes in the objective function over time (see  12). Increment-Delta-Bar-Delta (IDBD), introduced by  [34], has an adaptive mechanism based on the loss to adjust the learning rate $\eta_{i}$ for each sample $x_{i}$ for linear regression and has been extended to settings including RL [43]. Learning rate schedulers with time decay [31, 42] are coupled with optimizers, assuming gradual convergence to a good solution, but often require task-specific manual tuning. A meta-gradient reinforcement learning is proposed in [41], composed of a two-level optimization process: one that uses the agent’s objective and the other to learn meta-parameters of the objective function. Our work differs from these methods by employing a multi-armed bandit approach to dynamically select the learning rate over the training process, specifically targeting RL settings.
@@ 3 Preliminaries
This section introduces the Reinforcement Learning and Multi-Arm Bandits frameworks, defining supporting notation.
@@ 3.1 Deep Reinforcement Learning
An RL task is defined by a Markov Decision Process (MDP), that is by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see, e.g.35 for details). In RL, starting from an initial state $s_{0}$, a learner called agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$. In return, it receives a reward $r_{t}=R(s_{t},a_{t})$, reaching a new state $s_{t+1}$ according to the transition probability $P(s_{t},a_{t},\,\cdot\,)$. The agent’s objective is to learn a policy $\pi:\mathcal{S\times A}\rightarrow[0,1]$ which maps a distribution of actions given the current state, aiming to maximize expected returns $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl{[}\sum_{t=0}^{T}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\bigr{]}$.
To learn how to perform a task, value function-based algorithms coupled with ANNs [24, 25] approximate the quality of a given state-action pair $Q(s,a)$ using parameters $\theta$ to derive a policy $\pi_{\theta}(s)=\arg\max_{a\in\mathcal{A}}Q_{\theta}(s,a)$. By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ into the replay memory $\mathcal{D}$, the objective is to minimize the loss function $\mathcal{J}(\theta)$ defined by:
@@
where $\theta^{-}$ are the target network parameters used to compute the target of the learning network $y=r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})$. The parameters $\theta^{-}$ are periodically updated by copying the parameters $\theta$, leveraging stability during the learning process by fixing the target $y$. The minimization, and hence, the update of the parameters $\theta$, is done according to the optimizer’s routine. A simple possibility is to use stochastic gradient descent using mini-batch approximations of the loss gradient:
@@
$\theta_{n+1}\leftarrow\theta_{n}-\eta\nabla_{\theta}\mathcal{J}(\theta_{n})\,,\quad\text{where}\quad\nabla_{\theta}\mathcal{J}(\theta_{n})\approx\frac{1}{|\mathcal{B}|}\sum_{(s,a,r,s^{\prime})\in\mathcal{B}}2\bigl{(}Q_{\theta}(s,a)-y\bigr{)}\nabla_{\theta}Q_{\theta}(s,a)$ (2)
@@
with $\mathcal{B}$ being a mini-batch of transitions sampled from $\mathcal{D}$ and $\eta$ is a single scalar value called the learning rate. Unlike in supervised learning, where the loss function $\mathcal{J}(\theta)$ is typically stationary, RL presents a fundamentally different challenge: the policy is continuously evolving, leading to shifting distributions of states, actions, and rewards over time. This continuous evolution introduces instability during the learning process, which deep RL mitigates by employing a large replay memory and calculating the target using a frozen network with parameters $\theta^{-}$. However, stability also depends on how the parameters $\theta$ change during each update. This work aims to control these changes by dynamically selecting the learning rate $\eta$ over the training steps.
3.2 Multi-Armed Bandit
@@
Multi-Armed Bandits (MAB) provide an elegant framework for making sequential decisions under uncertainty (see for instance 20). MAB can be viewed as a special case of RL with a single state, where at each round $n$, the agent selects an arm $k_{n}\in\{1,\ldots,K\}$ from a set of $K$ arms and receives a feedback (reward) $f_{n}(k_{n})\in\mathbb{R}$. Like RL, MAB algorithms must balance the trade-off between exploring arms that have been tried less frequently and exploiting arms that have yielded higher rewards up to time $n$.
@@
To account for the non-stationarity of the RL rewards, we will consider in this work the MAB setting of adversarial bandits [6]. In this setting, at each round $n$, the agent selects an arm $k_{n}$ according to some distribution $p_{n}$ while the environment (the adversary) arbitrarily (e.g., without stationary constraints) determines the rewards $f_{n}(k)$ for all arms $k\in\mathcal{K}$. MAB algorithms are designed to minimize the pseudo-regret $G_{N}$ after $N$ rounds defined by:
$G_{N}=\max_{k\in\{1,\dots,K\}}\,\mathbb{E}\Biggl{[}\sum_{n=1}^{N}f_{n}(k)-\sum_{n=1}^{N}f_{n}(k_{n})\Biggr{]}\,,$
@@
where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}f_{n}(k)}$ represents the accumulated reward of the single best arm in hindsight, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm. A significant component in the adversarial setting is to ensure that each arm $k$ has a non-zero probability $p_{n}(k)>0$ of being selected at each round $n$: this guarantees exploration, which is essential for the algorithm’s robustness to environment changes.
@@ 4 Dynamic Learning Rate for Deep RL
In this section, we tackle the challenge of selecting the learning rate over the training steps by introducing a dynamic Learning Rate for deep Reinforcement Learning (LRRL). LRRL is a meta-learning approach designed to dynamically select the learning rate in response to the agent’s performance. LRRL couples with stochastic gradient descent optimizers and adapts the learning rate based on the reward achieved by the policy $\pi_{\theta}$ using an adversarial MAB algorithm. As the agent interacts with the environment, the average of observed rewards is used as bandit feedback to guide the selection of the most appropriate learning rate throughout the training process.
@@ 4.2 Updating the Probability Distribution
As we expect that the agent’s performance —and hence, the cumulative rewards— will improve over time, the MAB algorithm should receive non-stationary feedback. To take this non-stationary nature of the learning into account, we employ the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 6 for an introduction). At round $n$, Exp3 chooses the next arm (and its associated learning rate) according to the arm probability distribution $p_{n}$ which is based on weights $(w_{n}(k))_{1\leq k\leq K}$ updated recursively. Those weights incorporate a time-decay factor $\delta\in(0,1]$ that increases the importance of recent feedback, allowing the algorithm to respond more quickly to improvements in policy performance.
Specifically, after picking arm $k_{n}$ at round $n$, the RL agent interacts $C$ times with the environment and the MAB algorithm receives a feedback $f_{n}$ corresponding to the average reward of those $C$ interactions. Based on [26], this feedback is then used to compute the improvement in performance, denoted by $f^{\prime}_{n}$, obtained by subtracting the average of the past $j$ bandit feedbacks from the most recent one $f_{n}$:
@@
This update rule ensures that as the policy $\pi_{\theta}$ improves, the MAB algorithm continues to favor learning rates that are most beneficial under the current policy performance, thereby effectively handling the non-stationarity inherent in the learning process.

### v2

@@ Abstract
In deep Reinforcement Learning (RL) models trained using gradient-based techniques, the choice of optimizer and its learning rate are crucial to achieving good performance: higher learning rates can prevent the model from learning effectively, while lower ones might slow convergence. Additionally, due to the non-stationarity of the objective function, the best-performing learning rate can change over the training steps. To adapt the learning rate, a standard technique consists of using decay schedulers. However, these schedulers assume that the model is progressively approaching convergence, which may not always be true, leading to delayed or premature adjustments. In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a meta-learning approach that selects the learning rate based on the agent’s performance during training. LRRL is based on a multi-armed bandit algorithm, where each arm represents a different learning rate, and the bandit feedback is provided by the cumulative returns of the RL policy to update the arms’ probability distribution. Our empirical results demonstrate that LRRL can substantially improve the performance of deep RL algorithms for some tasks.
@@ 1 Introduction
Reinforcement Learning (RL), when combined with function approximators such as Artificial Neural Networks (ANNs), has shown success in learning policies that outperform humans in complex games by leveraging extensive datasets (see, e.g.37, 20, 43, 44). While ANNs were previously used as value function approximators [32], the introduction of Deep Q-Networks (DQN) by Mnih et al. [25, 26] marked a significant breakthrough by improving learning stability through two mechanisms: the target network and experience replay.
@@
Experience replay (see 23) stores the agent’s interactions within the environment, allowing sampling of past interactions in a random way that disrupts their correlation. The target network further stabilizes the learning process by periodically copying the parameters of the learning network. This strategy is necessary because the Bellman update —using estimations to update other estimations— would otherwise occur using the same network, potentially causing divergence. By leveraging the target network, gradient steps are directed towards a periodically fixed target, ensuring more stability in the learning process. Additionally, the learning rate [^1] The terms “learning rate” and “step-size” are used interchangeably in the literature and they technically refer to the same concept. hyperparameter controls the magnitude of these gradient steps in optimizers such as the Stochastic Gradient Descent (SGD) algorithm, affecting the training convergence.
@@
The learning rate is one of the most important hyperparameters, with previous work demonstrating that decreasing its value during policy finetuning can enhance performance by up to 25% in vanilla DQN [3]. Determining the appropriate learning rate is essential for achieving good model performance: higher values can prevent the agent from learning, while lower values can lead to slow convergence (see 15, 7, 46). However, finding a learning rate value that improves the model performance requires extensive and computationally expensive testing. In order to adapt its initial choice during training, optimizers such as Adam [18] and RMSProp [41] employ an internal scheme that dynamically adjusts the learning rate, considering, for instance, past gradient information. Therefore, various learning rate scheduling strategies can be combined with the optimizer to decrease the learning rate and improve the convergence over the training steps.
@@
Standard learning rate schedulers typically decrease the learning rate based on training progress using, e.g., linear or exponential decay strategies [35, 46]. Unlike supervised learning, RL usually involves generating data by trading off between exploration (discovery of new states) and exploitation (refining of the agent’s knowledge). As the policy improves, the data distribution encountered by the agent becomes more concentrated, but this evolution occurs at a different pace than the overall training progress. For instance, some environments require extensive exploration due to the sparseness of rewards, while others need more exploitation to refine the policy to the complexity of the task. Consequently, a more sophisticated decaying learning rate strategy that accounts for policy performance rather than training steps can significantly enhance learning in deep RL.
Contributions.
@@
In this work, we propose dynamic Learning Rate for deep Reinforcement Learning (LRRL), a method to select the learning rate on the fly for deep RL. Our approach acknowledges that different learning phases require different learning rates, and as such, instead of scheduling the learning rate decay using some blanket approach, we dynamically choose the learning rate using a Multi-Armed Bandit (MAB) algorithm, which accounts for the current policy’s performance. Our method has the advantage of being algorithm-agnostic and applicable to any optimizer. By exploiting different configurations, we aim to illustrate the use cases of LRRL and assess its robustness. We provide experiments with the Adam and RMSProp optimizers, with different sets of arms, and two deep RL algorithms (DQN and IQN). Tests are carried out across a variety of Atari games, using baselines provided in the Dopamine framework (see 9). We also show the effectiveness of LRRL using a standard SGD optimizer in a non-RL optimization context. Related work is detailed in Appendix A. Our main contributions and results can be summarized as follows:
  • We introduce LRRL, a novel approach that leverages a MAB algorithm to dynamically select the learning rate during training in deep RL.
  • We assess LRRL robustness across a set of Atari games, using different configurations, optimizers and deep RL algorithms. Our empirical results demonstrate that LRRL achieves performance competitive with or superior to standard learning using fixed learning rates or decay schedulers. We further highlight its advantages by considering an SGD optimizer on stationary non-convex landscapes.
  • More generally, we show that LRRL matches the scheduler’s behavior by dynamically choosing lower learning rate values as the training progresses.
@@ 2 Preliminaries
This section introduces the RL and MAB frameworks, defining supporting notation.
@@ 2.1 Deep Reinforcement Learning
An RL task is defined by a Markov Decision Process (MDP), that is by a tuple $\mathcal{(S,A},P,R,\gamma,T)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the set of possible actions, $P:\mathcal{S\times A\times S\rightarrow}[0,1]$ the transition probability, $R:\mathcal{S\times A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1]$ the discount factor, and $T$ the horizon length in episodic settings (see, e.g.39 for details). In RL, starting from an initial state $s_{0}$, a learner called the agent interacts with the environment by picking, at time $t$, an action $a_{t}$ depending on the current state $s_{t}$. In return, it receives a reward $r_{t}=R(s_{t},a_{t})$, reaching a new state $s_{t+1}$ according to the transition probability $P(\cdot|s_{t},a_{t})$. The agent’s objective is to learn a policy $\pi(\cdot|s)$ that maps a state $s$ to a probability distribution over actions in $\mathcal{A}$, aiming to maximize expected returns $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl{[}\sum_{t=0}^{T}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\bigr{]}$.
To learn how to perform a task, value function-based algorithms coupled with ANNs [25, 26] approximate the quality of a given state-action pair $Q(s,a)$ using parameters $\theta$ to derive a policy $\pi_{\theta}(\cdot|s)=1_{a^{*}(s)}(\cdot)$ where $1_{a^{*}(s)}$ denotes the uniform distribution over ${a^{*}(s)}$ with $a^{*}(s)=\arg\max_{a\in\mathcal{A}}Q_{\theta}(s,a)$. By storing transitions $(s,a,r,s^{\prime})=(s_{t},a_{t},r_{t},s_{t+1})$ into the replay memory $\mathcal{D}$, the objective is to minimize the loss function defined by:
@@
where $\theta^{-}$ are the target network parameters used to compute the target of the learning network $y=r+\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q_{{\theta}^{-}}(s^{\prime},a^{\prime})$. The parameters $\theta^{-}$ are periodically updated by copying the parameters $\theta$, leveraging stability during the learning process by fixing the target $y$. The minimization, and hence the update of the parameters $\theta$, is done according to the optimizer’s routine. A simple possibility is to use SGD using mini-batch approximations of the loss gradient:
@@
$\begin{gathered}\theta_{n+1}\leftarrow\theta_{n}-\eta\nabla_{\theta}\mathcal{J}(\theta_{n})\,,\text{where}\\ \nabla_{\theta}\mathcal{J}(\theta_{n})\approx\frac{1}{|\mathcal{B}|}\sum_{(s,a,r,s^{\prime})\in\mathcal{B}}2\bigl{(}Q_{\theta}(s,a)-y\bigr{)}\nabla_{\theta}Q_{\theta}(s,a),\end{gathered}$ (2)
@@
with $\mathcal{B}$ being a mini-batch of transitions sampled from $\mathcal{D}$ and $\eta$ is a single scalar value called the learning rate. Unlike in supervised learning, where the loss function $\mathcal{J}(\theta)$ is typically stationary, RL presents a fundamentally different challenge: the policy is continuously evolving, leading to shifting distributions of states, actions, and rewards over time. This continuous evolution introduces instability during learning, which deep RL mitigates by employing a large replay memory and calculating the target using a frozen network with parameters $\theta^{-}$. However, stability also depends on how the parameters $\theta$ change during each update. This work aims to adapt to these changes by dynamically selecting the learning rate $\eta$ over the training steps.
2.2 Multi-Armed Bandits
@@
MAB provide an elegant framework for making sequential decisions under uncertainty (see for instance 21). MAB can be viewed as a special case of RL with a single state, where at each round $n$, the agent selects an arm $k_{n}\in\{1,\ldots,K\}$ from a set of $K$ arms and receives a feedback (reward) $f_{n}(k_{n})\in\mathbb{R}$. Like RL, MAB algorithms must balance the trade-off between exploring arms that have been tried less frequently and exploiting arms that have yielded higher rewards up to time $n$.
@@
To account for the non-stationarity of RL rewards, we will consider in this work the MAB setting of adversarial bandits [6]. In this setting, at each round $n$, the agent selects an arm $k_{n}$ according to some distribution $p_{n}$ while the environment (the adversary) arbitrarily (e.g., without stationary constraints) determines the rewards $f_{n}(k)$ for all arms $k\in\mathcal{K}$. MAB algorithms are designed to minimize the pseudo-regret $G_{N}$ after $N$ rounds defined by:
$G_{N}=\mathbb{E}\Biggl{[}\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)-\sum_{n=1}^{N}f_{n}(k_{n})\Biggr{]}\,,$
@@
where the randomness of the expectation depends on the MAB algorithm and on the adversarial environment, $\smash{\sum_{n=1}^{N}\max_{k\in\{1,\dots,K\}}f_{n}(k)}$ represents the accumulated reward of the (pontentially changing) single best arm, and $\smash{\sum_{n=1}^{N}f_{n}(k_{n})}$ is the accumulated reward obtained by the algorithm. A significant component in the adversarial setting is to ensure that each arm $k$ has a non-zero probability $p_{n}(k)>0$ of being selected at each round $n$: this guarantees exploration, which is essential for the algorithm’s robustness to environment changes.
@@ 3 Dynamic Learning Rate for Deep RL
In this section, we tackle the challenge of selecting the learning rate over the training steps by introducing a dynamic LRRL. LRRL is a meta-learning approach designed to dynamically select the learning rate in response to the agent’s performance. LRRL couples with SGD optimizers and adapts the learning rate based on the reward achieved by the policy $\pi_{\theta}$ using an adversarial MAB algorithm. As the agent interacts with the environment, the average of observed rewards is used as bandit feedback to guide the selection of the most appropriate learning rate throughout the training process.
@@ 3.2 Updating the Probability Distribution
As we expect that the agent’s performance —and hence, the cumulative rewards— will improve over time, the MAB algorithm should receive non-stationary feedback. To take this non-stationary nature of the learning into account, we employ a modified version of the Exponential-weight algorithm for Exploration and Exploitation (Exp3, see 6 for an introduction). At round $n$, Exp3 chooses the next arm (and its associated learning rate) according to the arm probability distribution $p_{n}$ which is based on weights $(w_{n}(k))_{1\leq k\leq K}$ updated recursively. Those weights incorporate a time-decay factor $\delta\in(0,1]$ that increases the importance of recent feedback, allowing the algorithm to respond more quickly to changes of the best action and to improvements in policy performance.
Specifically, after picking arm $k_{n}$ at round $n$, the RL agent interacts $C$ times with the environment and the MAB algorithm receives a feedback $f_{n}$ corresponding to the average reward of those $C$ interactions. Based on Moskovitz et al. [27], this feedback is then used to compute the improvement in performance, denoted by $f^{\prime}_{n}$, obtained by subtracting the average of the past $j$ bandit feedbacks from the most recent one $f_{n}$:
@@
This update rule ensures that as the policy $\pi_{\theta}$ improves the cumulative reward, the MAB algorithm continues to favor learning rates that are most beneficial under the current policy performance, thereby effectively handling the non-stationarity inherent in the learning process.